lilaspie7:

Find C^2: 24^2 + 7^2 = c^2

Serenity1jackson:

um idk

lian:

c=25

lilaspie7:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Serenity1jackson um idk \(\color{#0cbb34}{\text{End of Quote}}\) If you dont know the answer, don't say anything

lilaspie7:

thx lian

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lian c=25 \(\color{#0cbb34}{\text{End of Quote}}\) direct answer

lian:

no problem

lilaspie7:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lian no problem \(\color{#0cbb34}{\text{End of Quote}}\) medal 4 me pls? lol

Eiwoh2:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lian c=25 \(\color{#0cbb34}{\text{End of Quote}}\) Would you like to explain how you reached that answer? Direct answers are a big no-no on here.

lilaspie7:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lian c=25 \(\color{#0cbb34}{\text{End of Quote}}\) break it down please on how you got 25

lian:

576+72=c2

lian:

576+49=c2

lilaspie7:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lian 576+49=c2 \(\color{#0cbb34}{\text{End of Quote}}\) lol

lian:

576+49=c2

lian:

625=c2

Eiwoh2:

Keep it in one message, please. Wouldn’t want to crowd the thread.

lian:

625-c2=0 -c2+625=0

lilaspie7:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lian 625-c2=0 -c2+625=0 \(\color{#0cbb34}{\text{End of Quote}}\) that doesnt seem right

lian:

well, thats not all of it

Eiwoh2:

I’d simply suggest not giving a direct answer next time. Our goal here is to teach and guide the user, remember. That’s all. :)

lian:

ok, sorry

Eiwoh2:

You’re fine, live and learn 👌

XioGonz:

\[((24^2)+72)=0\] \[24=2^3*3\] \[(24)^2 = (2^3•3)2 = 2^6 • 3^2\] \[((2^6•3^2) + 7^2) - c^2 = 0 \] Factor: 625-c^2 We know that a difference of two perfect squares, --> A2 - B2 can be factored into: (A+B) • (A-B) AB = BA = commutative property of multiplication. - AB + AB = zero so it can be eliminated from the expression. \[(c + 25) • (25 - c) = 0 \] \[ c+25 = 0 \] Subtract 25 from both sides \[ c = -25\] -c+25 = 0 Subtract 25 from both sides of the equation : \[ -c = -25\] Multiply both sides of the equation by (-1) : c = 25